OpenStudy (anonymous):

i need help with 3 problems that i've done on my own is anyone willing to help me?

OpenStudy (anonymous):

maybe :D what sort of help do you need?

OpenStudy (anonymous):

Two boys want to use a seesaw and they need to move the seesaw so that their weights will balance out. The formula is given by w1 • d1 = w2 • d2 where w1 = weight of the first boy, d1 = distance of the first boy from the fulcrum, w2 = weight of the second boy, and d2 = distance of the second boy from the fulcrum. Rewrite the formula to solve for d2. Show all steps in your work .w1*d1=w2*d2 w2*d2/w2 = (w1*d1)/w2 d2 = w1*d1/w2

OpenStudy (anonymous):

Okay... nothing wrong with it, and from the looks of it, that's already done.

OpenStudy (anonymous):

Marcie wants to enclose her yard with a fence. Her yard is in the shape of a triangle attached to a rectangle.The area of this figure can be found by the formula A = (wh) + 0.5(bh). If Marcie wants the total area to be larger than a specified value, she can use the formula A > (wh)+ 0.5(bh). Rewrite this formula to solve for b. Show all steps in your work.

OpenStudy (anonymous):

i need help on that. i'm confused.

OpenStudy (anonymous):

what's confusing? just solve for b as if that > sign were a = sign instead.

OpenStudy (anonymous):

can you show me in steps?

OpenStudy (anonymous):

sure :) \[\large A > wh+0.5\color{green}bh\] Now, bring everything without b to the left side. \[\large A - wh > 0.5 \color{green}b h\] Should be easy from here? :)

OpenStudy (anonymous):

A > (wh)+ 0.5(bh)

OpenStudy (anonymous):

2(a-wh)/h>b

OpenStudy (anonymous):

That's it :D

OpenStudy (anonymous):

This is the last one. Find a formula that is used in sports or leisure activities. Be sure that that formula has at least two variables. Rewrite your formula to solve for the other variable. Show all steps in your work. Bicycling-distance=rate*time Soccer-Distance=time*speed

OpenStudy (anonymous):

That's what i got so far.

OpenStudy (anonymous):

not sure about this one, sorry ':)

OpenStudy (anonymous):

i will post it. Thank you for all your help:)